# HaHeAE: Learning Generalisable Joint Representations of Human Hand and Head Movements in Extended Reality

### v1

@@ Abstract
Human hand and head movements are the most pervasive input modalities in extended reality (XR) and are significant for a wide range of applications. However, prior works on hand and head modelling in XR only explored a single modality or focused on specific applications. We present HaHeAE – a novel self-supervised method for learning generalisable joint representations of hand and head movements in XR. At the core of our method is an autoencoder (AE) that uses a graph convolutional network-based semantic encoder and a diffusion-based stochastic encoder to learn the joint semantic and stochastic representations of hand-head movements. It also features a diffusion-based decoder to reconstruct the original signals. Through extensive evaluations on three public XR datasets, we show that our method 1) significantly outperforms commonly used self-supervised methods by up to 74.0% in terms of reconstruction quality and is generalisable across users, activities, and XR environments, 2) enables new applications, including interpretable hand-head cluster identification and variable hand-head movement generation, and 3) can serve as an effective feature extractor for downstream tasks. Together, these results demonstrate the effectiveness of our method and underline the potential of self-supervised methods for jointly modelling hand-head behaviours in extended reality.
@@ I Introduction
Recently, self-supervised representation learning has emerged as a promising paradigm for learning latent semantic embeddings of speech [10, 11], mouse movements [12, 13], or gaze behaviours [14]. Self-supervised methods use the original input data as their own supervisions without requiring extra human annotations, which are usually costly, cumbersome, and time-consuming [15]. These methods can also be well generalised across different tasks [16, 17]. Despite the advantages and potential of self-supervised representation learning, no prior work has explored such approaches for learning joint representations of human hand and head behaviours.
To fill this gap, we introduce HaHeAE – the first self-supervised method to learn generalisable joint representations of human hand and head behaviours in extended reality. At the core of our method is an autoencoder (AE) that uses a graph convolutional network-based (GCN-based) semantic encoder to learn the joint semantic representation of hand-head movements and employs a diffusion-based stochastic encoder to encode the remaining stochastic variations. Then, a diffusion-based decoder is applied to reconstruct the original signals from the semantic and stochastic representations. In the training process, we propose to use hand-head movement forecasting as an auxiliary task to enhance the spatial-temporal features encoded in the semantic representation. We conduct extensive experiments on three publicly available XR datasets, i.e. the EgoBody [18], Aria digital twin (ADT) [19], and GIMO [20] datasets that contain human hand and head data collected during various daily activities from different users in diverse XR environments, and demonstrate that our method significantly outperforms the self-supervised methods that are commonly used for different signals by up to 74.0% in terms of reconstruction quality and is generalisable across users, activities, and XR environments. We further show that the semantic representation learned by our method can be used to identify hand-head behaviour clusters with human-interpretable semantics. In contrast, the stochastic representation can generate variable hand-head movements. We also evaluate the practical use of our method as a generic feature extractor for two practical sample downstream tasks, i.e. user identification and activity recognition, which are essential for intelligent XR systems to understand users and the interaction context [5, 21, 4, 22], and demonstrate consistent improvements on both tasks. The full source code and trained models will be released upon acceptance.
@@ II-A Hand and Head Behaviour Modelling
Understanding and modelling human behaviours is an essential research topic in extended reality [23, 24, 14, 7]. Human hand and head movements are particularly significant as they are the most pervasive input modalities in extended reality and are considered crucial components for future human-aware intelligent XR systems [5, 4, 3]. Against this background, researchers have devoted tremendous efforts to the computational modelling of human hand and head behaviours for a wide range of XR applications. Specifically, Belardinelli et al. modelled human hand behaviour using Gaussian hidden Markov models to predict users’ target objects in interactive virtual environments [1]. Bachynskyi et al. explored second- and third-order lag models for modelling human hand dynamics for mid-air pointing tasks in extended reality [25]. Gamage et al. proposed a hybrid classical-regressive model to learn hand kinematics for predicting continuous 3D hand trajectory for ballistic movements in immersive virtual environments [6]. Yang et al. employed a hierarchical Bayesian long short-term memory (LSTM) network to encode historical head movements to predict future head trajectory on omnidirectional images [26]. Hu et al. used a 1D convolutional neural network (CNN) to extract features from users’ head orientations to further predict human visual attention in immersive virtual environments [8, 21]. Hu et al. proposed to use a combination of a 1D CNN, a recurrent neural network (RNN), and a multi-layer perceptron (MLP) to model human head behaviours for the recognition of user activities in 360-degree videos [5]. However, existing methods for modelling human hand and head behaviours typically explored only a single modality and focused on a specific application. In stark contrast, we are the first to learn generalisable joint representations of human hand and head movements in extended reality.
@@ II-B Self-Supervised Representation Learning
Self-supervised representation learning is a machine learning paradigm where models are trained using only the original input data as supervision information without requiring additional human annotations. Self-supervised methods have recently been demonstrated to be highly effective for learning generalisable latent semantic embeddings of various signals including speech [11, 10], mouse movements [12, 13], and gaze behaviours [14]. Specifically, Rekimoto et al. used a speech-to-unit encoder to generate hidden speech units and employed a unit-to-speech decoder to reconstruct speech from the encoded speech units [11]. They demonstrated that the learned hidden speech units can be generalised to different speech types (whispered and normal speech) and different users. Zhang et al. used a Transformer-based encoder-decoder architecture to learn semantic representations of mouse behaviours from both continuous mouse cursor locations and discrete mouse events (click vs. movements) [12]. They showed that the learned mouse representations can be applied to identify interpretable mouse behaviour clusters, retrieve similar mouse trajectories, as well as benefit downstream tasks such as next activity prediction. Jiao et al. proposed an implicit neural representation learning-based method to learn semantic representations from low-resolution eye gaze data [14]. The learned representations can be used to up-sample the gaze data to arbitrary resolutions and improve performance on the downstream task of gaze-based user identification. Despite the potential of self-supervised representation learning, no prior work has explored such approaches for learning joint semantic representations of human hand and head behaviours.
@@ III-A Problem Definition
We formulate the problem of learning joint representations of human hand and head movements as a self-supervised task. This task involves learning latent embeddings from the original hand-head movements and using the learned representation to reconstruct the original signals. Given that different XR environments may use different coordinate systems, we propose to represent hand-head movements using relative coordinates with the origin set to the position of the head. This approach also yields a generalisable representation of hand-head data collected across different XR environments. We use the 3D positions of the left and right hands $ha\in R^{6}$ to represent the hand movement and employ head orientation $he\in R^{3}$ to denote the head movement, where $he$ is a unit vector indicating head forward direction. Given a sequence of hand movements $HA_{1:N}=\{ha_{1},ha_{2},...,ha_{N}\}\in R^{6\times N}$ and head orientations $HE_{1:N}=\{he_{1},he_{2},...,he_{N}\}\in R^{3\times N}$, the task is to learn latent embeddings that can reconstruct the original signals.
@@ III-B GCN-based Semantic Encoder
Inspired by the fact that graph convolutional networks surpass other architectures such as Transformers or RNNs in processing human pose data [34, 35], we proposed to use a GCN-based encoder to extract semantic features from the hand-head data. Specifically, we modelled the hand-head data $H_{1:N}=\{ha_{1},he_{1},ha_{2},he_{2},...,ha_{N},he_{N}\}\in R^{9\times N}$ as fully connected spatial and temporal graphs $H\in R^{3\times 3\times N}$ with their adjacency matrices measuring the weights between each pair of nodes. The spatial graph consists of three joints representing the head, left hand, and right hand, respectively, while the temporal graph contains $N$ nodes corresponding to hand-head data at different time steps.
We first used a spatial-temporal graph convolutional network (ST-GCN) to map the original hand-head data into a latent feature space. The ST-GCN first performed temporal convolution by multiplying the data with a temporal adjacency matrix $A^{T}\in R^{N\times N}$, then mapped the original node features ($3$ dimensions) into latent space ($16$ dimensions) using a feature matrix $W\in R^{3\times 16}$, and finally performed spatial convolution by multiplying the data with a spatial adjacency matrix $A^{S}\in R^{3\times 3}$ to obtain hand-head features $H\in R^{16\times 3\times N}$. We then used a residual GCN module containing two GCN blocks to further process the hand-head features. Each GCN block consists of an ST-GCN to extract features, a layer normalisation (LN) to normalise the data, a Tanh activation function, and a dropout layer with a dropout rate of $0.1$ to prevent the GCN from overfitting. The feature matrix of the ST-GCN used in the GCN block was set to $W\in R^{16\times 16}$ to ensure that the input and output of the block had the same size. A residual connection was applied for each GCN block to improve the network flow. We used three residual GCN modules in total to enhance the hand-head features with a downsample layer applied after the first two modules. The downsample layer used a 1D average pooling with a kernel size of two to compress the hand-head data along the temporal dimension. After the downsample operation, we obtained hand-head features in the size of $H\in R^{16\times 3\times(N/4)}$.
@@ IV-B Evaluation Settings
We set the sequence length of hand-head movements $N$ to $40$ and the time horizon of hand-head forecasting $\Delta n$ to $3$. We trained our method and the baseline methods using the Adam optimiser with a learning rate of $41e-$ 4. We used a batch size of $64$ to train these methods for a total of $130$ epochs. All the methods were implemented using the PyTorch framework.
@@ IV-C Reconstruction Results
We first trained and tested our method on the EgoBody dataset and indicated the reconstruction performances of different methods in Table I. The table shows the MPJPE error (in centimetres) of hand trajectory reconstruction and the mean angular error (in degrees) of head orientation reconstruction. As can be seen from the table, our method significantly outperforms other methods on EgoBody, achieving an improvement of $51.9\%$ ($1.66$ vs $3.45$) and $64.5\%$ ($0.83$ vs $2.34$) in terms of hand and head reconstruction, respectively. We further used a paired Wilcoxon signed-rank test to compare the performance of our method with that of the baseline methods. We validated that our method’s and the baselines’ differences are statistically significant ($p<0.01$).
TABLE I: Hand and head reconstruction errors (hand unit: centimetres, head unit: degrees) of different representation methods on the EgoBody, ADT, and GIMO datasets. All the methods are trained only on the EgoBody dataset and tested on all three datasets, with the best results marked in bold. Our method significantly outperforms other methods on all the datasets in terms of both hand and head reconstruction quality. $E_{sem}$ and $E_{sto}$ are essential for reconstructing the original signals.
@@
VAE_1DCNN 3.57 2.55 3.88 2.78 4.42 3.10
VAE_LSTM 7.22 5.41 7.78 9.97 8.81 8.11
VAE_GRU 6.78 4.39 7.35 6.16 8.37 6.37
VAE_MLP 3.45 2.34 3.93 2.73 4.31 2.93
Ours w/o $E_{sem}$ 56.80 93.39 56.00 93.04 56.93 91.81
Ours w/o $E_{sto}$ 10.62 11.25 11.88 12.86 11.15 12.05
Ours 1.66 0.83 1.97 0.71 2.40 1.25
@@
We further tested our method and the baseline methods, which were trained on EgoBody, directly on the ADT and GIMO datasets without any fine-tuning. The hand and head reconstruction performances of different methods are summarised in Table I. We can see from the table that our method consistently outperforms other methods by a large margin on both the ADT and GIMO datasets. Specifically, on the ADT dataset, our method outperforms other methods by $49.2\%$ ($1.97$ vs. $3.88$) and $74.0\%$ ($0.71$ vs. $2.73$) in terms of hand and head reconstruction, respectively. On the GIMO dataset, our method achieves an improvement of $44.3\%$ ($2.40$ vs. $4.31$) and $57.3\%$ ($1.25$ vs. $2.93$) in hand and head reconstruction, respectively. A paired Wilcoxon signed-rank test was further performed to compare the performances of our method with other methods and the results demonstrated that the differences between our method and other methods on both the ADT and GIMO datasets are statistically significant ($p<0.01$).
@@
We also evaluated the effectiveness of $E_{sem}$ and $E_{sto}$ in reconstructing the original hand-head signals. Specifically, we first replaced $E_{sem}$ and $E_{sto}$ with Gaussian noise respectively and then used them to reconstruct the hand-head data, following prior works on evaluating representations [13, 33]. We can see from Table I that the reconstruction performance deteriorates significantly after removing $E_{sem}$ or $E_{sto}$, validating that both $E_{sem}$ and $E_{sto}$ are of great significance for reconstructing the original hand-head data. We also find that $E_{sem}$ has a notably higher influence on the reconstruction quality than $E_{sto}$, demonstrating that $E_{sem}$ captures more critical information of the hand-head data.
IV-D Summary
The above results validate that our method significantly outperforms commonly used self-supervised representation learning methods in reconstructing both hand trajectories and head orientations. More importantly, after being trained on one dataset, our method can be directly used on other datasets with unseen users, activities, and XR scenarios without fine-tuning to achieve high-quality reconstruction. These results demonstrate our method’s generalisation capability across different users, activities, and XR environments.

### v3

@@ Abstract
Human hand and head movements are the most pervasive input modalities in extended reality (XR) and are significant for a wide range of applications. However, prior works on hand and head modelling in XR only explored a single modality or focused on specific applications. We present HaHeAE – a novel self-supervised method for learning generalisable joint representations of hand and head movements in XR. At the core of our method is an autoencoder (AE) that uses a graph convolutional network-based semantic encoder and a diffusion-based stochastic encoder to learn the joint semantic and stochastic representations of hand-head movements. It also features a diffusion-based decoder to reconstruct the original signals. Through extensive evaluations on three public XR datasets, we show that our method 1) significantly outperforms commonly used self-supervised methods by up to 74.1% in terms of reconstruction quality and is generalisable across users, activities, and XR environments, 2) enables new applications, including interpretable hand-head cluster identification and variable hand-head movement generation, and 3) can serve as an effective feature extractor for downstream tasks. Together, these results demonstrate the effectiveness of our method and underline the potential of self-supervised methods for jointly modelling hand-head behaviours in extended reality.
@@ I Introduction
Recently, self-supervised representation learning has emerged as a promising paradigm for learning latent semantic embeddings of speech [10, 11], mouse movements [12, 13], or gaze behaviours [14]. Self-supervised methods use the original input data as their own supervision without requiring extra human annotations, which are usually costly, cumbersome, and time-consuming [15]. These methods can also be well generalised across different tasks [16, 17]. Despite the advantages and potential of self-supervised representation learning, no prior work has explored such approaches for learning joint representations of human hand and head behaviours.
To fill this gap, we introduce HaHeAE – the first self-supervised method to learn generalisable joint representations of human hand and head behaviours in extended reality. At the core of our method is an autoencoder (AE) that uses a graph convolutional network-based (GCN-based) semantic encoder to learn the joint semantic representation of hand-head movements and employs a diffusion-based stochastic encoder to encode the remaining stochastic variations. Then, a diffusion-based decoder is applied to reconstruct the original signals from the semantic and stochastic representations. In the training process, we propose to use hand-head movement forecasting as an auxiliary task to enhance the spatial-temporal features encoded in the semantic representation. We conduct extensive experiments on three publicly available XR datasets, i.e. the EgoBody [18], Aria digital twin (ADT) [19], and GIMO [20] datasets that contain human hand and head data collected during various daily activities from different users in diverse XR environments, and demonstrate that our method significantly outperforms the self-supervised methods that are commonly used for different signals by up to 74.1% in terms of reconstruction quality and is generalisable across users, activities, and XR environments. We further show that the semantic representation learned by our method can be used to identify hand-head behaviour clusters with human-interpretable semantics. In contrast, the stochastic representation can generate variable hand-head movements. We also evaluate the practical use of our method as a generic feature extractor for two practical sample downstream tasks, i.e. user identification and activity recognition, which are essential for intelligent XR systems to understand users and the interaction context [5, 21, 4, 22], and demonstrate consistent improvements on both tasks. The full source code and trained models are available at https://git.hcics.simtech.uni-stuttgart.de/public-projects/HaHeAE.
@@ II-A Hand and Head Behaviour Modelling
Understanding and modelling human behaviours is an essential research topic in extended reality [23, 24, 14, 7]. Human hand and head movements are particularly significant as they are the most pervasive input modalities in extended reality and are considered crucial components for future human-aware intelligent XR systems [5, 4, 3]. Against this background, researchers have devoted tremendous efforts to the computational modelling of human hand and head behaviours for a wide range of XR applications. Specifically, Belardinelli et al. modelled human hand behaviour using Gaussian hidden Markov models to predict users’ target objects in interactive virtual environments [1]. Bachynskyi et al. explored second- and third-order lag models for modelling human hand dynamics for mid-air pointing tasks in extended reality [25]. Gamage et al. proposed a hybrid classical-regressive model to learn hand kinematics for predicting continuous 3D hand trajectory for ballistic movements in immersive virtual environments [6]. Yang et al. employed a hierarchical Bayesian long short-term memory (LSTM) network to encode historical head movements to predict future head trajectory on omnidirectional images [26]. Hu et al. used a 1D convolutional neural network (CNN) to extract features from users’ head orientations to further predict human visual attention in virtual environments [8, 21, 27]. Hu et al. proposed to use a combination of a 1D CNN, a recurrent neural network (RNN), and a multi-layer perceptron (MLP) to model human head behaviours for the recognition of user activities in 360-degree videos [5]. However, existing methods for modelling human hand and head behaviours typically explored only a single modality and focused on specific applications rather than learning generalisable representations. In stark contrast, we are the first to learn generalisable joint representations of human hand and head movements in extended reality.
@@ II-B Self-Supervised Representation Learning
Self-supervised representation learning is a machine learning paradigm where models are trained using only the original input data as supervision information without requiring additional human annotations. Self-supervised methods have recently been demonstrated to be highly effective for learning generalisable latent semantic embeddings of various signals including speech [11, 10], mouse movements [12, 13], and gaze behaviours [14]. Specifically, Rekimoto et al. used a speech-to-unit encoder to generate hidden speech units and employed a unit-to-speech decoder to reconstruct speech from the encoded speech units [11]. Zhang et al. used a Transformer-based encoder-decoder architecture to learn semantic representations of mouse behaviours from both continuous mouse cursor locations and discrete mouse events (click vs. movements) [12]. Jiao et al. proposed an implicit neural representation learning-based method to learn semantic representations from low-resolution eye gaze data [14]. Despite the potential of self-supervised representation learning, no prior work has explored such approaches for learning joint semantic representations of human hand and head behaviours.
@@ III-A Problem Definition
We formulate the problem of learning joint representations of hand and head movements as a self-supervised task. This task involves learning latent embeddings from the original hand-head movements and using the learned representation to reconstruct the original signals. Given that different XR environments may use different coordinate systems, we propose to represent hand-head movements using relative coordinates with the origin set to the position of the head. This approach yields a generalisable representation of hand-head data collected across different XR environments. We use 3D positions of the left and right hands $ha\in R^{6}$ to represent hand movement following prior works on hand behaviour modelling [1, 25, 6] and employ head orientation $he\in R^{3}$ to denote head movement, where $he$ is a unit vector indicating head forward direction. Given a sequence of hand movements $HA_{1:N}=\{ha_{1},ha_{2},...,ha_{N}\}\in R^{6\times N}$ and head orientations $HE_{1:N}=\{he_{1},he_{2},...,he_{N}\}\in R^{3\times N}$, the task is to learn latent embeddings that can reconstruct the original signals.
@@ III-B GCN-based Semantic Encoder
Inspired by the fact that graph convolutional networks surpass other architectures such as CNNs or RNNs in learning correlations of different body parts [35, 36, 37], we proposed to use a GCN-based encoder to extract semantic features from the hand-head data. Specifically, we modelled the hand-head data $H_{1:N}=\{ha_{1},he_{1},ha_{2},he_{2},...,ha_{N},he_{N}\}\in R^{9\times N}$ as fully connected spatial and temporal graphs $H\in R^{3\times 3\times N}$ with their adjacency matrices measuring the weights between each pair of nodes. The spatial graph consists of three joints representing the head, left hand, and right hand, respectively, while the temporal graph contains $N$ nodes corresponding to hand-head data at different time steps.
We first used a spatial-temporal graph convolutional network (ST-GCN) to map the original hand-head data into a latent feature space (see Figure 1). The ST-GCN first performed temporal convolution by multiplying the data with a temporal adjacency matrix $A^{T}\in R^{N\times N}$, then mapped the original node features ($3$ dimensions) into latent space ($16$ dimensions) using a feature matrix $W\in R^{3\times 16}$, and finally performed spatial convolution by multiplying the data with a spatial adjacency matrix $A^{S}\in R^{3\times 3}$ to obtain hand-head features $H\in R^{16\times 3\times N}$. We then used a residual GCN module containing two GCN blocks to further process the hand-head features. Each GCN block consists of an ST-GCN to extract features, a layer normalisation (LN) to normalise the data, a Tanh activation function, and a dropout layer with a dropout rate of $0.1$ to prevent the GCN from overfitting. The feature matrix of the ST-GCN used in the GCN block was set to $W\in R^{16\times 16}$ to ensure that the input and output of the block had the same size. A residual connection was applied for each GCN block to improve the network flow. We used three residual GCN modules in total to enhance the hand-head features with a downsample layer applied after the first two modules. The downsample layer used a 1D average pooling with a kernel size of two to compress the hand-head data along the temporal dimension. After the downsample operation, we obtained hand-head features in the size of $H\in R^{16\times 3\times(N/4)}$.
@@ IV-B Evaluation Settings
We set the sequence length of hand-head movements $N$ to $40$ and the time horizon of hand-head forecasting $\Delta n$ to $3$. We trained our method and the baseline methods for a total of $130$ epochs using the Adam optimiser with a learning rate of $41e-$ 4 and a batch size of $64$. Our method was implemented on an NVIDIA Tesla V100 SXM2 32GB GPU with an Intel(R) Xeon(R) Platinum 8260 CPU @ 2.40GHz and it took around nine hours for our method to finish its training on the EgoBody dataset.
@@ IV-C Reconstruction Results
We first trained and tested our method on the EgoBody dataset and indicated the reconstruction performances of different methods in Table I. The table shows the MPJPE error (in centimetres) of hand trajectory reconstruction and the mean angular error (in degrees) of head orientation reconstruction. As can be seen from the table, our method significantly outperforms other methods on EgoBody, achieving an improvement of $51.8\%$ ($1.664$ vs. $3.455$) and $64.3\%$ ($0.834$ vs. $2.338$) in terms of hand and head reconstruction, respectively. We further performed a paired Wilcoxon signed-rank test by first generating a distribution of performance metrics for each method and then comparing them directly. The test sample size of the EgoBody dataset is $66293$. The results validated that the differences between our method and the baselines are statistically significant ($p<0.01$). We also analysed the distributions of different methods’ reconstruction errors and validated that our method achieves better performance than other methods (see supplementary material for more details).
TABLE I: Hand and head reconstruction errors (hand unit: centimetres, head unit: degrees) of different representation methods on the EgoBody, ADT, and GIMO datasets. Best results are in bold.
@@
VAE_1DCNN 3.575 2.549 3.876 2.776 4.422 3.100
VAE_LSTM 7.254 5.421 7.933 9.928 8.908 8.060
VAE_GRU 6.776 4.390 7.351 6.161 8.369 6.371
VAE_MLP 3.455 2.338 3.932 2.733 4.310 2.927
Ours 1.664 0.834 1.966 0.707 2.397 1.247
Ours_1DCNN 2.010 1.070 2.370 1.095 2.883 1.500
Ours_LSTM 1.706 0.842 1.937 0.713 2.587 1.341
Ours_GRU 1.715 0.861 1.964 0.718 2.658 1.377
Ours_MLP 1.840 0.851 2.213 0.822 2.660 1.279
Ours w/o $E_{sem}$ 56.803 93.361 55.952 92.991 57.005 91.864
Ours w/o $E_{sto}$ 10.604 11.278 11.889 12.878 11.158 12.042
@@
We further tested our method and the baseline methods, which were trained on EgoBody, directly on the ADT and GIMO datasets without any fine-tuning. The hand and head reconstruction performances of different methods are summarised in Table I. We can see from the table that our method consistently outperforms other methods by a large margin on both the ADT and GIMO datasets. Specifically, on the ADT dataset, our method outperforms other methods by $49.3\%$ ($1.966$ vs. $3.876$) and $74.1\%$ ($0.707$ vs. $2.733$) in terms of hand and head reconstruction, respectively. On the GIMO dataset, our method achieves an improvement of $44.4\%$ ($2.397$ vs. $4.310$) and $57.4\%$ ($1.247$ vs. $2.927$) in hand and head reconstruction, respectively. A paired Wilcoxon signed-rank test was further performed to compare the performances of our method with other methods. The test sample sizes of the ADT and GIMO datasets are $91998$ and $44081$, respectively. The results demonstrated that the differences between our method and other methods on both the ADT and GIMO datasets are statistically significant ($p<0.01$).
The Effectiveness of the GCN-based Semantic Encoder
To test the effectiveness of our GCN-based encoder, we replaced it with the baseline VAE encoders (1DCNN, LSTM, GRU, or MLP) to re-train our method. We can see from Table I that our GCN-based encoder achieves better performance than other encoders in terms of hand-head reconstruction quality, demonstrating its superiority for encoding hand-head signals.
@@
We also evaluated the effectiveness of $E_{sem}$ and $E_{sto}$ in reconstructing the original signals. Specifically, we first replaced $E_{sem}$ and $E_{sto}$ with Gaussian noise respectively and then used them to reconstruct the hand-head data, following prior works on evaluating representations [13, 34]. We can see from Table I that the reconstruction performance deteriorates significantly after removing $E_{sem}$ or $E_{sto}$, validating that both $E_{sem}$ and $E_{sto}$ are significant for reconstructing the original hand-head data. We also find that $E_{sem}$ has a notably higher influence on the reconstruction quality than $E_{sto}$. This is because $E_{sem}$ captures the semantic information that is critical for reconstruction while $E_{sto}$ encodes the remaining stochastic variations that are less important [34, 13].